OpenStudy (anonymous):

Find the y-intercept for the line whose equation is given below. -4x + 2y = 10

OpenStudy (lgbasallote):

put it in the form y = mx + b first

OpenStudy (anonymous):

x = 0 2y = 10 y = 5 (0,5)

OpenStudy (anonymous):

Y=mx+c y intercept would be c

OpenStudy (anonymous):

TO find the yintercept replace x with zero if x is zero we have 0+2y= 10 we are looking for Y we divide both side by 2 since the coeficient of y is 2 Y=5